Joe and mbsool gradpoint.com/Frame/component/courseplayer?enrolmentid 12208613 GRADPOI 35. A spring is stretched from equnbrium to 0.90 m, and the work done on the spring equals 18 what is the value of the spring constant? 44 N/m 32 N/m 11 N/m start 11:32 AM 1206 home end 5€ 6
Answer #1

Work done to stretch the spring a distance x is,

Work done,W = 0.5kx^2

k = W/(0.5x^2) = 18J/(0.5*0.9^2)m^2 = 44.45 N.m